# Deep generative modelling of canonical ensemble with differentiable thermal properties

### v1

@@ Abstract
We propose a variational modelling method with differentiable temperature for canonical ensembles. Using a deep generative model, the free energy is estimated and minimized simultaneously in a continuous temperature range. At optimal, this generative model is a Boltzmann distribution with temperature dependence. The training process requires no dataset, and works with arbitrary explicit density generative models. We applied our method to study the phase transitions (PT) in the Ising and XY models, and showed that the direct-sampling simulation of our model is as accurate as the Markov Chain Monte Carlo (MCMC) simulation, but more efficient. Moreover, our method can give thermodynamic quantities as differentiable functions of temperature akin to an analytical solution. The free energy aligns closely with the exact one to the second-order derivative, so this inclusion of temperature dependence enables the otherwise biased variational model to capture the subtle thermal effects at the PTs. These findings shed light on the direct simulation of physical systems using deep generative models.
The canonical ensemble is for a system in thermal equilibrium with a heat bath, where the temperature, as a principal thermodynamic variable, determines the probability distribution of states. The key to calculating macroscopic thermal properties is the partition function:
[…]
where $E$ is the energy function, $\beta$ is the inverse of the temperature $T$, $F$ is the Helmholtz free energy. For example, after obtaining the partition function $\mathcal{Z}$, the mean energy and the heat capacity can be readily calculated as
@@
For most physical systems, it is impossible to analytically solve $\mathcal{Z}$: an exact enumeration belongs to the class of #P-hard [1]. In practical calculations, it is common to approximate $\mathcal{Z}$ numerically[2, 3]. However, even numerical approximation remains highly challenging, especially for systems undergoing phase transitions (PTs), where the partition function must account for a large number of metastable states.
Due to the complexity of $\mathcal{Z}$, sample-based methods are widely used in numerical calculations, among which the Markov Chain Monte Carlo (MCMC) method holds prominence. The MCMC method is an importance sampling technique giving the Boltzmann distribution in the canonical ensemble:
@@
It can provide the prevailing sample configurations at a microscopic level, which are often absent in partition function approaches. This microscopic information is very valuable, e.g., in the identification of intermediate states in chemical reactions. Macroscopic properties can also be estimated by averaging the samples [4], namely, statistical averaging. An inherent limitation of MCMC is its efficiency. Due to the Markov chain structure of the rejection sampling, it takes multiple timesteps to produce an independent and identically distributed (i.i.d.) sample, which prohibits the MCMC from performing direct sampling. [5]. This autocorrelation problem becomes even worse near PTs [5, 6]. Additionally, unlike the explicit temperature dependence in the partition function, the MCMC method typically requires more computational effort for results on multiple temperature points, further diminishing its overall efficiency.
Recently, the explicit density generative model, a deep learning method, was introduced to overcome the efficiency limitation of MCMC [7, 8, 9]. It can direct give i.i.d. samples from parameterized distributions to compute statistical averages [7]. Some of the models can also provide the free energy estimation [8, 10], which is deemed challenging for MCMC. However, the limitation of temperature dependence persists, as one still has to train separate models for each temperature point, similar to MCMC. What makes it worse than MCMC is that, during this discrete-temperature training, the functional dependence of free energy and other thermodynamic quantities on temperature is broken, a major trade-off due to the absence of a differentiable $\mathcal{Z}$. In the case of MCMC, despite its inefficiency due to discrete temperature, one can still rely on its unbiased estimation to get accurate results. However, deep learning models, being variational models, may exhibit biases and lack physical fidelity. Consequently, the accuracy of deep generative models is often inadequate for directly representing the ensemble, particularly during phase transitions. When applying these variational models, one usually has to compensate with the aid of the MCMC rejection [8] or multiple-temperature training like annealing [10].
Variational temperature-differentiable optimization.
@@
In this Letter, we propose a general framework of directly modelling the canonical ensemble in a continuous temperature range. The resulting deep generative model is a direct-sampling model of the target Boltzmann distribution with temperature dependence. As the model is differentiable with temperature, the estimates, including the free energy and the partition function, can be differentiated in a similar way to an analytical solution.
@@
First, considering a single temperature, one can minimize the reverse Kullback-Leibler divergence (KLD) [11] to optimize an explicit density generative model, i.e.,
[…]
where $q_{\theta}$ is the model distribution, and $p$ is the Boltzmann distribution (Eq. (3)) at this temperature. After optimization, the model $q_{\theta}$ is an approximation of $p$. Additional, because the $D_{KL}\geq 0$, the statistical average, $\mathbb{E}_{\mathbf{x}\sim q_{\theta}}[\log q_{\theta}(\mathbf{x})+\beta E(\mathbf{x})]$, has a variational lower bound at $-\log\mathcal{Z}$, it can be used as an estimate of the partition function in Eq. (1[8, 10]. This has been successfully demonstrated on various applications [12, 13, 14, 15, 16]. However, as the true Boltzmann distribution is a function of temperature, training schemes optimizing on a single temperature may be biased and unphysical. This is usually demonstrated as the hardness of optimizing Eq. (4).
Here, we propose that approximating the temperature-dependent Boltzmann distribution by training in a differentiable temperature range. Conventionally, finding the temperature-dependent Boltzmann distribution involves the following free energy minimization problem.
$\begin{split}\min_{q_{\theta}}\sum_{\{\mathbf{x}\}}\left[1/\beta\cdot q_{\theta}(\mathbf{x},\beta)\log q_{\theta}(\mathbf{x},\beta)+q_{\theta}(\mathbf{x},\beta)E(\mathbf{x})\right],\\ \text{s.t.}~{}\sum_{\{\mathbf{x}\}}q_{\theta}(\mathbf{x},\beta)=1,~{}\forall\beta,\end{split}$ (5)
One can see that the Boltzmann distribution in Eq. (3) is the solution [17].
Next, we use an explicit density model as the variational $q_{\theta}$, and convert Eq. (5) into a similar form as Eq. (4), so that we can perform deep-learning training. As explicit density models are normalized distributions, the normalization is naturally met. The summation over $\{\bm{x}\}$ becomes an estimation over $q_{\theta}$. And this minimum is also the minimum of an integration of the free energy. We further turn this integration into an estimation over an equal-probability distribution of $\beta$. Thus, the following loss function is equivalent to the free energy minimization in Eq. (5):
@@
$\mathcal{L}=\mathop{\mathbb{E}}_{\begin{subarray}{c}\beta\sim\text{U}\\ \mathbf{x}\sim q_{\theta}(\cdot,\beta)\end{subarray}}\left[\log q_{\theta}(\mathbf{x},\beta)+\beta E(\mathbf{x})\right]=\mathop{\mathbb{E}}_{\beta\sim\text{U}}\left[-\log\bar{\mathcal{Z}}(\beta)\right].$ (6)
@@
One can perform the stochastic gradient descent (SGD) [18, 19] to optimize it. Gradients can be obtained from the computational graph and automatic differentiation, which are widely used in deep learning [20]. After training, we obtain $q_{\theta}$ as a variational approximation of the Boltzmann distribution with differentiable temperature dependence.
@@
Compared to Eq. (4), this variational temperature-differentiable (VaTD) loss function is an integration of the reverse KLD over a continuous temperature range. Using the reverse KLD as loss, this training doesn’t require any dataset, because the sample $\mathbf{x}$ is drawn from a direct-sample generative model $q_{\theta}$. Another benefit is that one gets a temperature-differentiable estimate of the free energy: $\log\bar{\mathcal{Z}}(\beta)$, so its derivatives with respect to temperature, e.g., the mean energy in Eq. (2) and heat capacity in Eq. (2), can be readily calculated using automatic differentiation. Other thermodynamic quantities, e.g., the magnetization, can be also estimated by computing the statistical averages of directly sampled batches, and they are also functions of differentiable temperature.
Figure 1: (a) The exact and estimated free energy of the 2D Ising model on a $16\times 16$ PBC square lattice, and the relative error between the two (inset). The temperature factor (-$T$) is removed for better comparison. (b) The estimated square magnetization of the Ising model. (c) Configurations directly sampled from the learned model at three temperatures, with background color as a continuous interpolation of the discrete site for better visualization. (d) The mean energy and heat capacity of the Ising model, estimated using the differentiation of the free energy, compared with the results obtained from the MCMC simulation.
One concern is that $q_{\theta}$ may not be flexible enough to cover the whole space of probability distribution, which may result in sub-optimal solutions for Eq. (5) and miss the Boltzmann distribution. However, this problem can be overcome by employing sufficiently large neural network models to representing $q_{\theta}$, as guaranteed by the universal approximation theorem [21]. Additionally, compared with directly constructing a specific model always following the Boltzmann distribution [22], our method comes with nearly no restriction on the model, and thus has maximum fitting ability. In the VaTD training, the model initially starts as a random non-Boltzmann distribution, and is optimized to be a Boltzmann distribution, which extends the generality of the proposed framework. One can choose specially tailored models for different applications.
@@
In performing SGD, one should notice that the gradients of Eq. (6) is not trivial. As the model parameters $\theta$ are used in the expectation operator, the nice form of statistical averaging will be lost in the gradients. To overcome this, two numerical methods, namely reparameterization and reinforce estimation, have been proposed [23]. They enable the swapping of the derivative operation with the expectation operation, thereby converting the derivative calculations of statistical averages into the statistical averaging of derivatives. The same problem reoccurs when calculating the derivative of estimated thermal differentiable function with respect to $\beta$. This can be solved again by using these two methods. Typically only the first-order forms of these two methods are presented for purpose of training. We extend them to the second order in appendix [17]. It is worth noting that higher-order generalizations are feasible as long as the neural network remains continuous.
Numerical experiments.
In the numerical experiments, we demonstrate our method using two famous statistical models: two-dimensional Ising and XY models. The former one uses a discrete-variable model, PixelCNN [24], while the latter one uses a continuous-variable model, the normalizing flow (NF).
[…]
For Ising, we chose the PixelCNN model [24], which is an explicit density generative model for discrete-variable distributions. Its probability distribution is decomposed into the product of a series of conditional probability distributions, i.e., $q_{\theta}(\mathbf{x},\beta)=\prod_{i}q_{\theta}(x_{i}|x_{0},\cdots,x_{i-1};\beta)$. The conditional distribution we chose is the Bernoulli distribution, $x_{i}\sim B(1,p)$, where $p$ is the probability of a spin-up state, obtained from a ResNet [25] given the preceding variables $\{x_{0},\cdots,x_{i-1}\}$ and $\beta$. To have a unified input form, we repeated $\beta$ to match the size of $\{x_{0},\cdots,x_{i-1}\}$, and then concatenated them along the channel dimension. The PixelCNN model generates variables sequentially, one at a time. During each step, a variable is stochastically sampled from the Bernoulli distribution. The conditional causality is established through the dependence of the Bernoulli parameter $p$ on the preceding variables. The model was trained in a temperature range of $\beta\in[0.05,1.2]$. To eliminate the $\mathbb{Z}_{2}$ symmetry, we fixed the first spin to be spin-up.
After training, the PixelCNN model can perform direct sampling, and the sampled batches were used for statistical averaging. In Fig. 1(a), we estimated the free energy as a continuous function of temperature using Eq. (6). As the 2D Ising can be analytically solved [26], we compared our numerical results with the exact free energy at each temperature point in Fig. 1(a), and found that the relative errors are small, with the maximum happening near the PT region where the system exhibits long-range correlations and multiple meta-stable states. Similarly, we also calculated the square magnetization, which is in excellent agreement with the result obtained by the MCMC simulation, as shown in Fig. 1(b). This indicates our method has comparable accuracy with MCMC, but is more efficient. The sampled configurations also provide valuable insights into the microscopic behaviors of the system, and demonstrates that the underlying physical transition is successfully captured. In Fig. 1(c), we observe that the trained model effectively captures the microscopic changes at lower $\beta$, where the uniform spin direction gradually collapses. Hence, our method trained a neural network model that undergoes the same microscopic thermal transition as the target Ising model.
To further prove that our model successfully captures the subtle thermal transitions near the PTs, we computed mean energy and heat capacity, which are the derivatives of the free energy, as shown in Eq. (2), and compared with the unbiased MCMC results in Fig. 1(d). Overall, our results agree very well with those from the MCMC simulation, with only slight deviations observed near the PT region. This implies that the estimated free energy aligns closely with the exact one to the second-order derivative. Moverover, comparing with statistical averaging using direct sampling, these derivatives have advantage in efficiency and accuracy (see appendix [17] for the detailed comparisons).
@@
Figure 2: (a) The estimated free energy and its standard deviation (inset) of the 2D XY model on a $16\times 16$ PBC square lattice. The temperature factor (-$T$) is removed for better comparison. (b) The estimated square magnetization of the XY model. (c) Configurations directly sampled from the learned model at three temperatures, the background color represents the vorticity. (d) The mean energy and heat capacity of the XY model, estimated using the differentiation of the free energy, compared with the results obtained from the MCMC simulation.
@@
The XY model is more complex than the Ising model. As an example of the BKT transition [27, 28], the XY model finds practical applications in explaining various systems, including thin ${}^{4}\mathrm{He}$ films [29], and quasi-2D layered superconductors [30]. The energy function of the XY model is $E(\mathbf{x})=-\sum_{<i,j>}\cos(x_{i}-x_{j})$, where $x_{i}$ and $x_{j}$ are angles in the continuous range of $[-\pi,\pi]$. Similar to the Ising case, we also used a 2D $16\times 16$ square lattice with PBC for the XY model.
@@
We chose NF models for continuous variables in a closed range [31, 32, 33, 34, 35]. In a NF model, the samples are initially drawn from a prior distribution, which is easy to sample from. We chose the von Mises distribution ,$\text{VM}(\mu,\kappa)$ where $\mu$ controls the mean and $\kappa$ controls the variance [36, 37], as the prior distribution of angles. To have a temperature-differentiable prior distribution, we modified $\kappa$ as $\kappa_{\beta}=\kappa\cdot\beta$. As VMs are sampled by an acceptance-rejection method, a special reparameterization should be used to optimize the $\mu$ and $\kappa$ [38, 17] (see the appendix). The sample $\bm{z}$ from the VM prior distribution was then transformed using parameterized invertible transformations, i.e., $\bm{x}=f_{\theta}(\bm{z},\beta)$. To incorporate temperature, we again repeated $\beta$ and padded it with the sample variables in the channel dimension, which is similar to the Ising case. The effects of these invertible transformations, besides changing the sample variables, also introduce probability changes to the VM prior distribution. These distribution changes are measured by the determinant of the Jacobian matrices [31]. The invertible transformation we chose is the piece-wise cubic spline transformation [35]. This kind of transformation is designed for continuous variables in a closed range. Then, the final probability of a sample is the product of the prior probability and the determinant of the Jacobian, $q_{\theta}(\mathbf{x},\beta)=\text{VM}(\bm{z}|\mu,\kappa_{\beta})\cdot|\text{det}\frac{\partial\bm{z}}{\partial\mathbf{x}}|$. This model was trained in the temperature range of $\beta\in[0.4,2.0]$. The $O(1)$ symmetry was eliminated by pinning the first spin to 00.
@@
After training, we performed direct sampling and used the obtained batches to perform statistical averaging. Fig. 2(a) shows the estimated free energy as a continuous function of temperature. As XY is not exactly solvable, we calculated the standard deviation (STD) to evaluate the estimation quality. According to Eq. (4), a perfect model would produce uniform estimates regardless of $\bm{x}$, so the small STDs in Fig. 2(a) indicate very good simulation results. Fig. 2(b) shows the estimated square magnetization, in excellent agreement with the MCMC result. To study the microscopic behavior, we plotted the sampled configurations at different temperatures in Fig. 2(c), which clearly shows the microscopic changes occurring near the PT. Specifically, with decreasing temperature, the vortex density decreases, and a quasi long-range order emerges. This demonstrates a good capture of the underlying physical transition.

### v2

# Deep generative modeling of the canonical ensemble with differentiable thermal properties

@@ Abstract
It is a long-standing challenge to accurately and efficiently compute thermodynamic quantities of many-body systems at thermal equilibrium. The conventional methods, e.g., Markov chain Monte Carlo, require many steps to equilibrate. The recently developed deep learning methods can perform direct sampling, but only work at a single trained temperature point and risk biased sampling. Here, we propose a variational method for canonical ensembles with differentiable temperature, which gives thermodynamic quantities as continuous functions of temperature akin to an analytical solution. The proposed method is a general framework that works with any tractable density generative model. At optimal, the model is theoretically guaranteed to be the unbiased Boltzmann distribution. We validated our method by calculating phase transitions in the Ising and XY models, demonstrating that our direct-sampling simulations are as accurate as Markov chain Monte Carlo, but more efficient. Moreover, our differentiable free energy aligns closely with the exact one to the second-order derivative, indicating that the variational model captures the subtle thermal transitions at the phase transitions. This functional dependence on external parameters is a fundamental advancement in combining the exceptional fitting ability of deep learning with rigorous physical analysis.
The canonical ensemble is for a system in thermal equilibrium with a heat bath, where the temperature, as a principal thermodynamic variable, determines the probability distribution of states. Calculating canonical ensemble averages gives thermodynamic properties of condensed-matter many-body systems, e.g., interacting magnetic spins and solvated proteins. This calculation is closely related to the partition function:
$\mathcal{Z}=\int\exp(-\beta E(\mathbf{x}))d\mathbf{x},~F=-T\log\mathcal{Z},$ (1)
where $E$ is the energy function, $\beta$ is the inverse of the temperature $T$, and $F$ is the Helmholtz free energy. For example, after obtaining $\mathcal{Z}$, the mean energy and the heat capacity can be readily calculated as derivatives of the free energy,
@@
For most physical systems, it is impossible to analytically solve $\mathcal{Z}$: an exact enumeration belongs to the class of #P-hard [partitionFn]. In practical calculations, it is common to approximate $\mathcal{Z}$ numerically [trg, meanfield]. However, even numerical approximation remains highly challenging, especially for systems undergoing phase transitions (PTs), where the partition function must account for a large number of metastable states.
Because of the complexity of $\mathcal{Z}$, sample-based methods are widely used in numerical calculations, among which the Markov chain Monte Carlo (MCMC) method holds prominence. The MCMC method is an importance sampling technique giving the Boltzmann distribution of the canonical ensemble,
@@
It can provide the prevailing state configurations at a microscopic level, which are often unexplored in partition function approaches. This microscopic information is very valuable, e.g., in the identification of intermediate states in chemical reactions. Macroscopic properties, such as thermodynamic quantities, can also be estimated by averaging the samples [SMtextbook], namely, statistical averaging. As MCMC has the unbiased estimation guarantee, this statistical averaging is the ensemble averaging. An inherent limitation of MCMC is its efficiency. Because of the Markov chain structure of the rejection sampling, it takes multiple timesteps to equilibrate and to produce independent and identically distributed (i.i.d.) samples, which prohibits the MCMC from performing direct sampling [autocorrelation1]. This autocorrelation problem becomes even worse near PTs [autocorrelation1, autocorrelation2]. Additionally, unlike the explicit temperature dependence in the partition function, the MCMC method typically requires more computational effort for results at multiple temperature points, further diminishing its overall efficiency.
Recently, a type of deep neural network model was introduced to address the efficiency limitations of MCMC: the tractable density generative model [explicitDensity, prl, frankNoe]. As a generative model, it can directly produce one-shot i.i.d. samples from parameterized distributions, without requiring rejections. Additionally, as an tractable density, it provides the exact probability of the sample [explicitDensity]. Some of the models can also provide free energy estimation [prl, wuDianPRL], which is deemed challenging for MCMC. However, the limitation of temperature dependence persists, as separate models must be trained for each temperature point, forming an inefficient scheme similar to MCMC. Beyond the efficiency issue, this single-temperature training breaks the continuous temperature dependence of Boltzmann distribution. This is reflected in biases, making the sampled distribution favor a few high-probability states and ignoring the vast meta-stable states, a phenomenon also known as the model collapse. In the case of MCMC, despite inefficiency, its unbiased estimation promises accurate results. However, deep learning models, being variational, may exhibit biases and lack physical fidelity. When applying these models, one usually has to compensate with the aid of the MCMC rejection [prl, PREfromRef], annealing training [wuDianPRL], or sophisticated neural network architectures [PRE2fromRef]. As proposed below, biases can be mitigated by employing a temperature-dependent model that accounts for thermal effects in the distribution.
Variational temperature-differentiable method
@@
Here, we introduce the variational temperature-differentiable (VaTD) method as a general framework for modelling the canonical ensemble in a continuous temperature range; by a minimal adaption of incorporating temperature as an input parameter, one can turn any tractable density model into a VaTD model. After training with randomly sampled temperatures, this VaTD model can be an accurate direct-sampling model of the target Boltzmann distribution with temperature dependence. Moreover, as the model is differentiable with temperature, the estimates, including the free energy and the partition function, can be differentiated similarly to an analytical solution.
@@
First, considering a single temperature, one can minimize the reverse Kullback-Leibler divergence (KLD) [KLD] to optimize an tractable density generative model, i.e.,
@@
$D_{KL}(q_{\theta}||p)=\mathop{\mathbb{E}}_{\mathbf{x}\sim q_{\theta}}\log\frac{q_{\theta}(\mathbf{x})}{p(\mathbf{x})}=\log\mathcal{Z}+\mathop{\mathbb{E}}_{\mathbf{x}\sim q_{\theta}}\left[\log q_{\theta}(\mathbf{x})+\beta E(\mathbf{x})\right],$ (4)
where $q_{\theta}$ is the model distribution, and $p$ is the Boltzmann distribution at this fixed temperature (Eq. (3)). After optimization, the model $q_{\theta}$ is an approximation of $p$. Additional, because the $D_{KL}\geq 0$, the statistical average, $\mathbb{E}_{\mathbf{x}\sim q_{\theta}}[\log q_{\theta}(\mathbf{x})+\beta E(\mathbf{x})]$, has a variational lower bound at $\log\mathcal{Z}$, it can thus be used as an estimate of the partition function in Eq. (1[prl, wuDianPRL]. This has been successfully demonstrated in various applications [equivariantFlowLattice, lattice2, freePart1, freePart2, lw2023PRL]. However, since the actual Boltzmann distribution is temperature-dependent, training schemes that optimize for a single temperature may become biased toward the prevailing states at that temperature. Additionally, as shown in Eq. (4), the model is optimized based on its own samples, which can lead to persistent bias.
In the VaTD training, we propose to approximate the Boltzmann distribution by training in a continuous temperature range. Varying temperatures can reveal hidden local minima, a concept akin to parallel tempering in MCMC, where high temperatures are applied to surmount energy barriers. This motivates us to derive a temperature-dependent loss function from the variational definition of the Boltzmann distribution, i.e., the distribution minimizes free energy across a continuous temperature range:
$\begin{split}\min_{q_{\theta}}\sum_{\{\mathbf{x}\}}\left[1/\beta\cdot q_{\theta}(\mathbf{x},\beta)\log q_{\theta}(\mathbf{x},\beta)+q_{\theta}(\mathbf{x},\beta)E(\mathbf{x})\right],\\ \text{s.t.}~\sum_{\{\mathbf{x}\}}q_{\theta}(\mathbf{x},\beta)=1,~\forall\beta.\end{split}$ (5)
Using the Lagrange multiplier, one can see that the Boltzmann distribution in Eq. (3) is the solution [SM].
Having temperature as an input parameter, the VaTD models can be used as the variational $q_{\theta}(\bm{x},\beta)$. This allows Eq. (5) to be formulated in a form that resembles Eq. (4), enabling straightforward deep-learning training. As tractable density models, VaTD models are naturally normalized. We then turn the summation into estimation, and the minimum at each $\beta$ into the minimum of an estimation over uniformly distributed $\beta$. Thus, the following loss function is equivalent to the free energy minimization in Eq. (5):
$\mathcal{L}=\mathop{\mathbb{E}}_{\begin{subarray}{c}\beta\sim\text{U}\\ \mathbf{x}\sim q_{\theta}(\bm{x},\beta)\end{subarray}}\left[\log q_{\theta}(\mathbf{x},\beta)+\beta E(\mathbf{x})\right]=\mathop{\mathbb{E}}_{\beta\sim\text{U}}\left[-\log\bar{\mathcal{Z}}(\beta)\right].$ (6)
@@
The stochastic gradient descent (SGD) [SGD1, SGD2] can be performed to optimize it. Gradients can be obtained from the computational graph and automatic differentiation, which are widely used in deep learning [autodiff]. As ensured in Eq. (5), optimizing this loss drives the model toward the unbiased Boltzmann distribution. Hence, after training, the model can accurately learn the actual Boltzmann distribution along with its temperature dependence.
@@
Besides its thermodynamic origin, this loss function also represents an integration of the reverse KLD from Eq. (4) over a continuous temperature range. Consequently, this training doesn’t require any dataset, as the sample is drawn from a generative model by direct sampling. Another benefit is that one gets a temperature-differentiable estimate of the free energy, $\log\bar{\mathcal{Z}}(\beta)$. So, its derivatives with respect to temperature, e.g., the mean energy and heat capacity in Eq. (2), can be readily calculated using automatic differentiation. As shown in the SI [SM], this is more time efficient than the sample-based approach. Other thermodynamic quantities, e.g., magnetization, can be estimated by statistical averaging of directly sampled batches, and they are also differentiable functions of temperature.
One concern is that $q_{\theta}$ may not be flexible enough to cover the whole space of probability distribution, which may result in sub-optimal solutions for Eq. (5) and miss the Boltzmann distribution. However, as guaranteed by the universal approximation theorem [universalnet], a sufficiently large neural network model can represent any distribution. Since VaTD training is applicable to any tractable density model, it is easy to apply more advanced model structures. The VaTD training does not require a redesign of the model. Instead, it simply integrates temperature as an input, thus preserving the model’s fitting capabilities. In contrast to previous works that involve constructing specialized models designed to consistently follow the Boltzmann distribution (e.g., [NoeTemperatureFlow]), our method serves as a general training framework for any tractable density model. Additionally, in the VaTD training, the model begins as a random distribution and is subsequently optimized to approximate the unbiased Boltzmann distribution
@@
In performing SGD, it is worth noting that the gradients of Eq. (6) are not trivial. As the model parameters $\theta$ are used in the expectation operator, the nice form of statistical averaging will be lost in the gradients. To overcome this, two numerical methods, namely reparameterization and REINFORCE estimation, have been proposed [reparamAndreinforce]. They enable the swapping of the differentiation operation with the expectation operation, thereby converting the derivative calculations of statistical averages into the statistical averaging of derivatives. The same problem reoccurs when calculating the derivative of the estimated thermal differentiable function. This can be solved again by using these two methods. Typically, only the first-order forms of these two methods are presented for training. We extend them to the second order in SI [SM]. It is worth noting that higher-order generalizations are feasible as long as the neural network remains continuous.
@@
Figure 1: (a) The exact and estimated free energy of the 2D Ising model on a $16\times 16$ square lattice with PBC, and the relative error between the two (inset). The temperature factor (T-T) is removed from the free energy($T\log\mathcal{Z}$) for better comparison. (b) The estimated squared magnetization of the Ising model. (c) State configurations directly sampled from the trained model at three temperatures, with background color as a continuous interpolation of the discrete site for better visualization.
Figure 2: The mean energy and heat capacity of the Ising model, estimated using the differentiation of the free energy, compared with the results obtained from the MCMC simulation.
Numerical experiments
In the numerical experiments, we validated our method using two famous statistical models: two-dimensional Ising and XY models. The former one uses a discrete-variable model, PixelCNN [pixelRCNN], as the tractable density model used in the VaTD method, while the latter one uses a continuous-variable model, the normalizing flow (NF).
Our 2D Ising model is on a $16\times 16$ square lattice with periodic boundary conditions (PBCs). The energy function is given by $E(\mathbf{x})=-\sum_{<i,j>}x_{i}x_{j}$, where <i,j><i,j> means the two spins are nearest neighbors on the lattice. Variables of Ising are discrete, with each variable $x_{i}$ taking a value of either $1$ or 1-1, representing spin-up or spin-down, respectively.
For Ising, we incorporated the temperature as an input into the PixelCNN [pixelRCNN] as the VaTD model. PixelCNN is an autoregressive generative model for discrete-variable distributions. Our VaTD PixelCNN’s probability distribution is decomposed into the product of a series of conditional probability distributions, i.e., $q_{\theta}(\mathbf{x},\beta)=\prod_{i}q_{\theta}(x_{i}|x_{0},\cdots,x_{i-1};\beta)$. The conditional distribution we chose is the Bernoulli distribution, $x_{i}\sim B(1,p)$, where $p$ is the probability of $x_{i}$ being in the spin-up state, as obtained from a residual network [resnet] given the preceding variables $\{x_{0},\cdots,x_{i-1}\}$ and the temperature parameter $\beta$. To have a unified input form for the residual network, we repeated $\beta$ to match the size of $\{x_{0},\cdots,x_{i-1}\}$, and then concatenated them along the channel dimension. The PixelCNN model generates variables sequentially, one at a time. During each step, a variable is stochastically sampled from the Bernoulli distribution. The conditional causality is established through the dependence of the Bernoulli parameter $p$ on the preceding variables. The model was trained in a temperature range of $\beta\in[0.05,1.2]$. To eliminate the $\mathbb{Z}_{2}$ symmetry, we fixed the first spin to be spin-up.
After training, the VaTD PixelCNN model can perform direct sampling, and the sampled batches were used for statistical averaging. In Fig. 1(a), we estimated the free energy as a continuous function of temperature using Eq. (6). As the 2D Ising can be analytically solved [isingExact], we compared our numerical results with the exact free energy at each temperature point in Fig. 1(a), and found that the relative errors are small, with the maximum happening near the PT region where the system exhibits long-range correlations. Similarly, we also calculated the squared magnetization, which is in excellent agreement with the result obtained by the unbiased MCMC simulation, as shown in Fig. 1(b). This indicates that our statistical averages from direct sampling follow the actual ensemble averages closely, similar to MCMC but more efficient. The sampled configurations also provide valuable insights into the microscopic behaviors of the system, and demonstrate that the underlying physical transition is successfully captured. In Fig. 1(c), we observe that, at lower $\beta$, the uniform spin direction gradually collapses. Hence, our trained neural network model undergoes the same microscopic thermal transition as the target Ising model.
To further demonstrate that our model unbiasedly captures the subtle thermal transition effects near the PTs, we computed the mean energy and heat capacity, which are derivatives of the free energy as shown in Eq. (2), using automatic differentiation on the estimated free energy. We then compared with the unbiased MCMC results in Fig. 2. Overall, our results agree very well with those from the MCMC simulation, with only slight deviations observed near the PT region. This implies that the estimated free energy aligns closely with the exact one to the second-order derivative. Moreover, compared with statistical averaging using direct sampling, this free energy derivative estimation of mean energy and heat capacity has advantages in efficiency and accuracy (see SI [SM] for detailed comparisons).
Figure 3: (a) The estimated free energy and its standard deviation (inset) of the 2D XY model on a $16\times 16$ square lattice with PBC. The temperature factor (T-T) is removed from the free energy($T\log\mathcal{Z}$) for better comparison. (b) The estimated squared magnetization of the XY model. (c) State configurations directly sampled from the trained model at three temperatures, the background color represents the vorticity.
Figure 4: The mean energy and heat capacity of the XY model, estimated using the differentiation of the free energy, compared with the results obtained from the MCMC simulation.
@@
Next, the XY model is more complex than the Ising model. As an example of the Berezinskii-Kosterlitz- Thouless transition [xyAndKT1, xyAndKT2], the XY model finds practical applications in explaining various systems, including thin $^{4}\mathrm{He}$ films [XYhe], and quasi-2D layered superconductors [XYsc]. The energy function of the XY model is $E(\mathbf{x})=-\sum_{<i,j>}\cos(x_{i}-x_{j})$, where $x_{i}$ and $x_{j}$ are angles in the continuous range of $\pi,\pi]$. Similar to the Ising case, we also used a 2D $16\times 16$ square lattice with PBC.
@@
For XY’s continuous variables, we incorporated the temperature as an input into a NF model [normalizingflow1, normalizingflow2, pwlinearAndquadratic, nsf, cubic] as the VaTD model. In this VaTD NF model, the samples were initially drawn from a prior distribution, which is easy to sample from. We chose the von Mises distribution, $\text{VM}(\mu,\kappa)$ where $\mu$ controls the mean and $\kappa$ controls the variance [vonMisesSample2, vonMisesSample1], as the prior distribution. To have a temperature-dependent prior distribution, we modified $\kappa$ as $\kappa_{\beta}=\kappa\cdot\beta$. As VMs were sampled by an acceptance-rejection method, a special reparameterization should be used to optimize $\mu$ and $\kappa$ [acptReparam, SM] (see the SI). The sample $\bm{z}$ from the VM prior distribution was then transformed using parameterized invertible transformations, i.e., $\bm{x}=f_{\theta}(\bm{z},\beta)$. To incorporate temperature, we again repeated $\beta$ and padded it with the sample variables $\bm{z}$ in the channel dimension, which is similar to the Ising case. The effects of these invertible transformations, besides changing the sample variables, also introduce probability changes to the VM prior distribution. These distribution changes are measured by the determinant of the Jacobian matrix [normalizingflow1]. The invertible transformation we chose is the piece-wise cubic spline transformation [cubic]. This kind of transformation is designed for continuous variables in a closed range. Then, the final probability of a sample is the product of the prior probability and the determinant of the Jacobian, i.e., $q_{\theta}(\mathbf{x},\beta)=\text{VM}(\bm{z}|\mu,\kappa_{\beta})\cdot|\text{det}\frac{\partial\bm{z}}{\partial\mathbf{x}}|$ where $\bm{x}=f_{\theta}(\bm{z},\beta)$. This model was trained in the temperature range of $\beta\in[0.4,2.0]$. The $U(1)$ symmetry was eliminated by pinning the first spin to 0.
@@
After training, we performed direct sampling and used the obtained batches to perform statistical averaging. Fig. 3(a) shows the estimated free energy as a continuous function of temperature. As XY is not exactly solvable, we calculated the standard deviation (STD) to evaluate the estimation quality. According to Eq. (3), a perfectly trained model would produce uniform free energy estimates regardless of $\bm{x}$, so the small STDs in Fig. 3(a) indicate accurate alignment with the target Boltzmann distribution. Then, to demonstrate our statistical averages accurately follow the ensemble averages, we compared our direct-sampling simulation results with the unbiased MCMC results; Fig. 3(b) shows the estimated squared magnetization, in excellent agreement with the MCMC results. To study the microscopic behavior, we plotted the sampled configurations at different temperatures in Fig. 3(c), which clearly shows the microscopic changes occurring near the PT. Specifically, with decreasing temperature, the vortex density decreases, and a quasi-long-range order emerges. This demonstrates an accurate capture of the underlying physical transition.